# On the Capacity of Self-Attention

### v1

@@ 5 The power of multiple heads
Single head with compression. If one head serves all items, then $B=m$. Applying (2) gives the requirement $d_{k}\;\geq\;c_{2}\,\frac{m^{2}}{d_{\text{model}}^{2}}\,\log m,$ and since $h=1$ here, the total key dimension is $D_{K}=d_{k}$.
Multiple heads with compression. Construction II partitions the items into $\frac{m}{d_{\text{model}}}$ heads with $B=d_{\text{model}}$ per head. Plugging $B=d_{\text{model}}$ into (1) yields $N_{3}(B)=\Theta(\sqrt{d_{k}\log m})$. Taking $d_{k}=c_{3}\log m$ with $c_{3}$ larger than the constant in (1) ensures $N_{3}\leq c_{1}d_{k}$ w.h.p., and the total key dimension is $D_{K}\;=\;h\,d_{k}\;=\;O\!\Big(\frac{m}{d_{\text{model}}}\log m\Big).$
Consequence. As a result, in the compressive regime, if $m=\omega(d_{\text{model}})$, the single head requirement above implies asymptotically larger $D_{k}$ than the multihead construction. Or, equivalently, for a fixed $D_{k}$ budget, multiple heads can handle more edges (relationships) than a single head, even in a permutation graph. Multiple heads do not boost per‑head expressivity; they localize de‑embedding noise by reducing block size $B$, so that each head aggregates leakage over only fewer coordinates, bringing the noise to a manageable level. Note that this is not a lower bound for all conceivable single‑head designs, but it shows that within the de‑embedding to signature template we use, a single head cannot perform as well as multiple heads.
@@ 6 Experiments
We conduct experiments in an idealized self-attention setting, mirroring our theoretical model, to test several predictions. First, we compare the empirical minimum total key dimension, $\hat{D}_{K}^{\star}$, to the predicted theoretical scaling law of $\Theta\left(\frac{m\log m}{d_{\text{model}}}\right)$, noting that optimization may fail to find a solution matching the theoretical constructive bound. And second, we test predictions regarding head count: whether a multi-head advantage appears in permutation graphs and how the empirically optimal number of heads tracks the theory.
@@ 6.1 Results and comparison to theory
We probe capacity on permutation graphs with $m\!\in\!\{64,128,256,512\}$ and $d_{\text{model}}\!\in\!\{16,32,64\}$. For each $(m,d_{\text{model}})$ we sweep head counts $h\!\in\!\{1,2,4,8,16,32,64\}$ and several total key sizes $D_{K}{=}h\,d_{k}$ (multiple $D_{K}$ per $h$). Each configuration is trained from $10$ seeds with the protocol described above (AdamW, fixed embeddings, single global threshold $\tau$). We evaluate average test micro–F1 on a fixed held‑out set and define the empirical threshold
$D_{K}^{\star}\;=\;\min\{\,D_{K}\;:\;\exists h\ \text{s.t. mean test micro‑F1}\geq 0.99\,\}.$
We denote by $h^{\star}$ a head count that attains $D_{K}^{\star}$. Full grids and per‑config step limits are in App. A.2.
To isolate sequence‑length effects at fixed embedding compression, we also traverse the diagonal $r\!\stackrel{{\scriptstyle\text{def}}}{{=}}\!m/d_{\text{model}}{=}8$ with $(m,d_{\text{model}})\in\{(128,16),(256,32),(512,64),(1024,128),(2048,256),(4096,512)\}$, using $3$ seeds for the largest points and increased budgets (App. A.2). Finally, to probe extreme compression we include a second r=32r{=}32 point $(1024,32)$ (in addition to $(512,16)$). We also repeat a subset of runs with a variable context length (App. A.3).
@@
We observe: (i) a sharp F1 transition in a narrow $D_{K}$ window (capacity threshold) across all $(m,d_{\text{model}},h)$ (Fig. 2 below and Fig. 6 in App. A.2); (ii) a pronounced multi‑head advantage for many $(m,d_{\text{model}})$, even though each query has a single target—splitting a fixed $D_{K}$ across more heads reduces interference from superposition (Fig. 2); and (iii) the optimal head count increases with compression $r{=}m/d_{\text{model}}$ (Fig. 2), while per‑head width at the threshold remains modest.
[…]
Figure 2: More heads are needed as $m$ grows and as $d_{\text{model}}$ shrinks. See App. A.2 for error bar description.
@@
The minimum $D_{K}^{\star}$ grows rapidly with $m$ and decreases rapidly with $d_{\text{model}}$; exact values appear in Figure 5 (in the Appendix). A single head often fails to reach $0.99$ F1 within the scanned $D_{K}$ (e.g., $(m,d_{\text{model}})\!\in\!\{(512,64),(256,32)\}$, Fig. 6), whereas several small heads pass at substantially smaller $D_{K}$.
@@
At $D_{K}^{\star}$, per‑head widths are small: $d_{k}^{\star}\!\in\![5,24]$ on the base grid (median $11$), indicating gains come primarily from adding heads rather than making each head wide (Table 1; App. A.2).
Figure 3: Comparison of $D_{K}^{\star}$ to theoretical scaling law. $x$-axis: scaling law prediction. $y$-axis: observed behavior. See App. A.2 for error bar description.
@@
Fixed‑compression diagonal (r=8r{=}8).
Holding $r$ constant collapses the prediction to $D_{K}^{\star}\propto r\log m$, so the dependence on $m$ should be logarithmic. Along $(128,16)\!\to\!(4096,512)$ we observe roughly this behavior from $m{\geq}512$ onward (Fig. 4): $D_{K}^{\star}$ grows slowly while $m$ grows exponentially, matching the $\log m$ factor. The first two points are slightly conservative (smaller $d_{\text{model}}$) and align with the same $d_{\text{model}}\!\gtrsim\!\log m$ finite‑size effect. We also expect the optimal head count to be proportional to $r$; the observed results align well with this expectation.
@@
Empirical thresholds align closely with the $m\log m/d_{\text{model}}$ capacity law and expose a clear multi‑head advantage even for one‑target graphs. Discrepancies appear exactly where theory anticipates stronger superposition (small $d_{\text{model}}$ and very large $m$). Overall, allocating key–query budget across more heads with modest width is the efficient path to capacity in compressed embeddings. In Appendix A.3, provide evidence that these results are largely independent of context length. Specifically, thresholds were stable across $\ell\in\{16,32\}$ with a small shift only when testing at longer $\ell$ than used for training.
7 Lower Bound on Relational Graph Recognition
In this section, we provide our lower bound on any self‑attention mechanism that uniformly recovers every directed graph on $m$ items with exactly $m^{\prime}$ edges in our model from Section 3 under a fixed positive margin $\gamma$. The number of such graphs is $\binom{m(m-1)}{m^{\prime}}$; essentially what we show is that the QK parameters must carry (at least) the description length of the edge set, and thus the total key dimension
$D_{K}\;=\;\Omega\!\Big(\tfrac{\log\binom{m(m-1)}{m^{\prime}}}{d_{\text{model}}}\Big)=\Omega\Big(\tfrac{m^{\prime}\log(m^{2}/m^{\prime})}{d_{\text{model}}}\Big).$
The result is independent of parameter precision and applies to any context length $\ell\geq 2$.
We start with some preliminaries. We first point out that we can focus only on length‑2 contexts. Uniform correctness for RGR requires that, for every ordered pair $(u,v)\in V\times V$, the decision $(u,v)\in E$?” is the same in every context containing $u$ and $v$. In particular it must be correct in the length‑2 context $\mathcal{C}=(u,v)$. Hence any lower bound proved using only length‑2 contexts applies to the full problem. We next provide two structural reductions.
(i) Multi‑head to a single bilinear form. For a length‑2 context and any monotone per‑pair head aggregator (max, log‑sum‑exp, sum), replacing it by the sum only makes the model more permissive. Writing
$M\;=\;\sum_{k=1}^{h}W_{Q}^{(k)}W_{K}^{(k)\top}\ \in\ \mathbb{R}^{d_{\text{model}}\times d_{\text{model}}},$
the score on $(u,v)$ is the single bilinear form
$S(u,v)\;=\;\mathbf{x}_{u}^{\top}M\,\mathbf{x}_{v},$ (3)
and an edge is recognized iff $S(u,v)>\tau$ for a global threshold $\tau$. Since each summand $W_{Q}^{(k)}W_{K}^{(k)\top}$ has rank at most $d_{k}$,
$\mathrm{rank}(M)\ \leq\ \sum_{k=1}^{h}d_{k}\;=\;D_{K}.$ (4)
@@ 7 Lower Bound on Relational Graph Recognition
(ii) Edge‑set description length. Let $m=|V|$ and N=m(m1)N=m(m-1) be the number of ordered, loop‑free pairs. For any target size $m^{\prime}$, the family $\mathcal{G}_{m,m^{\prime}}$ of directed graphs with exactly $m^{\prime}$ edges has cardinality $\binom{N}{m^{\prime}}$; hence any procedure that can realize an arbitrary $E\subseteq[N]$ with $|E|=m^{\prime}$ must (implicitly) transmit at least
$L(m,m^{\prime})\;=\;\ln\binom{N}{m^{\prime}}$ (5)

### v3

@@ 5 The power of multiple heads
Single head with compression. If one head serves all items, then $B=m$ and $D_{K}=d_{k}$. Applying (2) implies
$D_{K}=\Omega\left(\frac{m^{2}}{d_{\text{model}}^{2}}\,\log m\right).$
Multiple heads with compression. Construction II partitions the items into $\frac{m}{d_{\text{model}}}$ heads with $B=d_{\text{model}}$ per head. Plugging $B=d_{\text{model}}$ into (1) yields $N_{3}(B)=\Theta(\sqrt{d_{k}\log m})$. Taking $d_{k}=c_{3}\log m$ with $c_{3}$ larger than the constant in (1) ensures $N_{3}\leq c_{1}d_{k}$ w.h.p., and the total key dimension is
$D_{K}\;=\;h\,d_{k}\;=\;O\!\Big(\frac{m}{d_{\text{model}}}\log m\Big).$
Consequence. The additional $\frac{m}{d_{\text{model}}}$ term for a single head implies that if $m=\omega(d_{\text{model}})$ (the compressive regime), the single head requirement above implies asymptotically larger $D_{k}$ than the multihead construction. Or, equivalently, for a fixed $D_{k}$ budget, multiple heads can handle more edges (relationships) than a single head, even in a permutation graph. Multiple heads do not boost per‑head expressivity; they localize de‑embedding noise by reducing block size $B$, so that each head aggregates leakage over fewer coordinates, bringing the noise to a manageable level. Note that this is not a lower bound for all conceivable single‑head designs, but it shows that within the de‑embedding to signature template we use, a single head cannot perform as well as multiple heads.
@@ 6 Experiments
We conduct experiments in a setting mirroring our theoretical model, to test several predictions. First, we compare the empirical minimum total key dimension, $\hat{D}_{K}^{\star}$, to the predicted theoretical scaling law of $\Theta\left(\frac{m\log m}{d_{\text{model}}}\right)$, noting that optimization may fail to find a solution matching the theoretical constructive bound. And second, we test predictions regarding head count: whether permutation graph exhibit a multi-head advantage and how the empirically optimal number of heads tracks the theory. Permutation graphs ensure each query has at most one correct in-context target, so improvements from multiple heads cannot be attributed to attending to multiple true neighbors.
@@ 6.1 Results and comparison to theory
We probe capacity on permutation graphs with $m\!\in\!\{64,128,256,512\}$ and $d_{\text{model}}\!\in\!\{16,32,64\}$. For each $(m,d_{\text{model}})$ we sweep head counts $h\!\in\!\{1,2,4,8,16,32,64\}$ and several total key sizes $D_{K}{=}h\,d_{k}$ (multiple $D_{K}$ per $h$). Each configuration is trained from $10$ seeds with the protocol described above (AdamW, fixed embeddings, single global threshold $\tau$). We evaluate average test micro–F1 on a fixed held‑out set and define the empirical threshold $D_{K}^{\star}\;=\;\min\{\,D_{K}\;:\;\exists h\ \text{s.t. mean test micro‑F1}\geq 0.99\,\}.$ We denote by $h^{\star}$ a head count that attains $D_{K}^{\star}$. Full grids and per‑config step limits are in App. A.2.
To isolate sequence‑length effects at fixed embedding compression, we also traverse the diagonal $r\!\stackrel{{\scriptstyle\text{def}}}{{=}}\!m/d_{\text{model}}{=}8$ with $(m,d_{\text{model}})\in\{(128,16),(256,32),(512,64),$ (1024,128),(2048,256),(4096,512)}(1024,128),(2048,256),(4096,512)\}, using $3$ seeds for the largest points and increased budgets (App. A.2). Finally, to probe extreme compression we include a second r=32r{=}32 point $(1024,32)$ (in addition to $(512,16)$).
[…]
Figure 2: $\mathbf{h^{*}}$ minimizing $\mathbf{D_{K}}$. More heads are needed as $m$ grows or $d_{\text{model}}$ shrinks. See App. A.2 for error bar description.
@@
We observe: (i) a sharp F1 transition in a narrow $D_{K}$ window (capacity threshold) across all $(m,d_{\text{model}},h)$ (Fig. 1 in this section and Fig. 9 in App. A.2); (ii) a pronounced multi‑head advantage for many $(m,d_{\text{model}})$, even though each query has a single target—splitting a fixed $D_{K}$ across more heads reduces interference from superposition (Fig. 2); and (iii) the optimal head count increases with compression $\lambda{=}m/d_{\text{model}}$ (Fig. 2), while per‑head width at the threshold is modest.
@@
The minimum $D_{K}^{\star}$ grows rapidly with $m$ and decreases rapidly with $d_{\text{model}}$; exact values appear in Figure 8 (App. A.2). A single head often fails to reach $0.99$ F1 within the scanned $D_{K}$ (e.g., $(m,d_{\text{model}})\!\in\!\{(512,64),(256,32)\}$, Fig. 9), whereas several small heads pass at substantially smaller $D_{K}$.
@@
At $D_{K}^{\star}$, per‑head widths are small: $d_{k}^{\star}\!\in\![5,24]$ on the base grid (median $11$), indicating gains come from adding heads rather than making each head wide (Table 1; App. A.2).
Figure 3: Empirical Validation of Capacity Law. The minimum key dimension $D_{K}^{\star}$ tracks the theoretical prediction $\frac{m\log m}{d_{\text{model}}}$ tightly ($R^{2}=0.944$). See App. A.2 for error bar description.
@@
Fixed‑compression diagonal ($\lambda{=}8$).
Holding $\lambda$ constant collapses the prediction to $D_{K}^{\star}\propto\lambda\log m$, so the dependence on $m$ should be logarithmic. Along $(128,16)\!\to\!(4096,512)$ we observe roughly this behavior from $m{\geq}512$ onward (Fig. 4): $D_{K}^{\star}$ grows slowly while $m$ grows exponentially, matching the $\log m$ factor. The first two points are slightly conservative (smaller $d_{\text{model}}$) and align with the same $d_{\text{model}}\!\gtrsim\!\log m$ finite‑size effect. We also expect the optimal head count to be proportional to $\lambda$; the observed results align well with this expectation.
The Appendix describes further experimental results. For denser (regular) graphs, we show that $D_{K}^{\star}$ and optimal head count scale as predicted by the theory: with the number of edges, demonstrating that edges not vertices define the constraint on capacity (App. A.3).
@@
Empirical thresholds align closely with the $m\log m/d_{\text{model}}$ capacity law and expose a clear multi‑head advantage even for one‑target graphs. Discrepancies appear exactly where theory anticipates stronger superposition (small $d_{\text{model}}$ and very large $m$). Overall, allocating key–query budget across more heads with modest width is the efficient path to capacity in compressed embeddings.
7 Robustness to Model Extensions
We here stress-test our findings from the previous section in progressively more realistic model variants, and see that even with these extensions, the same core phenomena predicted by the simpler analysis continue to govern performance. (1) We incorporate scaled-softmax (without an OV channel) and demonstrate that the multi-head advantage remains. This model version preserves the sharp transition but shifts the required $D_{K}$ to the right, consistent with the extra log factor in our softmax construction. (2) We add a value channel and train on message retrieval; this also yields a multi-head advantage, showing the effect is not an artifact of thresholded edge classification. (3) In a full single-layer Transformer block trained with frozen GPT-2 embeddings on an induction-style retrieval task that only requires attention to a single location, we again see a clear multi-head advantage.
7.1 Experiments incorporating softmax
We also conducted experiments with the softmax version of the idealized model. These experiments mirror our experiments in the base model, with only the following changes:
  • Scores are subject to a softmax along the context dimension.
  • Scores are scaled by $\sqrt{d_{k}}$.
  • Scores are aggregated over heads via a sum (instead of max).
  • Training was allowed to run for up to 80,000 steps.
We depict several examples of the resulting F1-$D_{K}$ curves in Figure 5. As in the model with max over heads (and no softmax), we here see a distinct multi-head advantage when in the compressed embedding range of $m\gg d_{\text{model}}$. We also see that the required $D_{K}$ to reach an F1 of 0.99 is shifted to the right from the experiments run in our core model - consistent with the additional $\log m$ factor in our constructions.
@@ 7.1 Experiments incorporating softmax
7.2 Extension to a value channel
We next study the empirical impact of adding a value channel to our model. All aspects of the data generation and optimization protocol remain in our core model (permutation graphs on $m$ nodes, frozen normalized embeddings $x_{i}\in\mathbb{R}^{d_{\text{model}}}$, contexts of length $\ell$ with target-in-context rate $\rho$, and AdamW with the same learning rate and regularization), except that we equip the layer with a value pathway and change the learning objective from edge classification to message retrieval.
Concretely, the key–query path is unchanged: we retain the same learned projections $W_{Q},W_{K}\in\mathbb{R}^{d_{\text{model}}\times D_{K}}$, partitioned into $h$ heads with per-head width $d_{k}=D_{K}/h$, and we continue to aggregate scores by an elementwise max over heads $S_{\max}=\max_{k}Q^{(k)}(K^{(k)})^{\top}$ without $1/\sqrt{d_{k}}$ scaling or softmax. The only architectural addition is a random value map $W_{V}\in\mathbb{R}^{d_{\text{model}}\times d_{\text{msg}}}$, drawn once at initialization and then frozen. Each node $i$ carries a “message” $y_{i}=x_{i}W_{V}\in\mathbb{R}^{d_{\text{msg}}}$ that lies in the span of the original embeddings, matching the theoretical assumption of a fixed random value channel.
Given a context $\mathcal{C}$ with embedding matrix $X_{\mathcal{C}}$, we reuse the same attention scores $S_{\max}\in\mathbb{R}^{\ell\times\ell}$ to mix value messages. Let $V_{\mathcal{C}}=X_{\mathcal{C}}W_{V}$ collect the in-context messages. For each position $i$ whose outgoing permutation neighbor $\pi(i)$ also appears in the context, we form a predicted neighbor message
$\hat{y}_{i}=\big(S_{\max}V_{\mathcal{C}}\big)_{i}$ (3)
using the raw (un-normalized) scores in $S_{\max}$ as mixing weights, and we supervise it toward the true neighbor message $y_{\pi(i)}=x_{\pi(i)}W_{V}$. Training minimizes the mean squared error between $\hat{y}_{i}$ and $y_{\pi(i)}$ over all such “valid” positions in the batch, i.e.,
$\mathcal{L}_{\text{MSE}}\;=\;\frac{1}{|\mathcal{I}_{\text{valid}}|}\sum_{i\in\mathcal{I}_{\text{valid}}}\big\|\hat{y}_{i}-y_{\pi(i)}\big\|_{2}^{2},$ (4)
where $\mathcal{I}_{\text{valid}}$ is the set of indices whose permutation neighbor lies in the same context. Early stopping now operate on validation MSE for this message-retrieval task (with the same check interval and patience as before), and we report test-set MSE as the primary metric. Micro-F1 and score margins induced by the learned $W_{Q},W_{K}$ no longer play any role in optimization or stopping.
Findings (Value Retrieval).
We observe that the multi-head advantage found with edge classification transfers directly to the message-retrieval task. Figure 6 displays the test MSE as a function of total key dimension $D_{K}$ for $m=64$ with embedding dimensions $d_{\text{model}}=16$ and $d_{\text{model}}=32$.
Multi-head advantage in compressed regimes.
In the compressed regime where $d_{\text{model}}\ll m$ (Figure 6, left, $d_{\text{model}}=16$), we observe a significant separation between single-head and multi-head performance. The single-head model ($h=1$) fails to reduce MSE significantly even as $D_{K}$ increases, plateauing at a high error rate. In contrast, models with $h\geq 4$ are able to leverage the key budget effectively, driving the MSE down sharply. This confirms that even when the task is soft message retrieval rather than hard binary classification, the geometric bottleneck of identifying the correct neighbor requires the noise suppression impact of multiple heads.
Figure 6: Value retrieval MSE mirrors capacity transitions. Mean test MSE vs. total key dimension $D_{K}$ for a value-retrieval task with frozen random value mappings (m=64m{=}64). Top ($d_{\text{model}}{=}16$): In the compressed regime, single-head attention (blue) fails to retrieve the correct message, while multi-head attention succeeds. Bottom ($d_{\text{model}}{=}32$): With less compression pressure, the gap narrows, though multi-head models still attain the lowest final MSE. Error bars are 95% CIs over 3 runs
Behavior in less compressed regimes.
When the embedding dimension is relaxed to $d_{\text{model}}=32$ (Figure 6, right), the necessity for multiple heads diminishes, consistent with our theoretical predictions. Here, a single head ($h=1$) is competitive, and in some low-$D_{K}$ settings even outperforms highly fragmented architectures (e.g., $h=16$, where $d_{k}$ becomes very small). However, intermediate head counts ($h=4,8$) still achieve the lowest ultimate MSE.
Takeaway.
These results demonstrate that our core findings are not an artifact of the thresholded classification objective. The attention mechanism’s ability to route information—specifically, to select the correct value vector to mix—is governed by the same geometric capacity constraints derived for the edge-existence problem.
7.3 Full transformer layer with frozen GPT-2 embeddings
We next turn to a much more complex and realistic scenario: a full single layer Transformer block used to solve a controlled retrieval problem similar to induction-head pointer indrection. We use a fairly standard structure: pre-LN attention plus MLP block with residuals, and use frozen GPT-2 token embeddings with a tied unembedding.
Frozen vocabulary and embedding space.
Let $W_{\mathrm{TE}}\in\mathbb{R}^{|\mathcal{V}|\times d_{\mathrm{model}}}$ be GPT-2-small’s token embedding matrix ($d_{\mathrm{model}}=768$). We restrict to an item set of size $m$, using the $m$ tokens with the highest unigram frequency in the WikiText-103 training corpus (estimated from the first 10 million tokens using the GPT-2 tokenizer), excluding the EOS token. This yields a frozen table
$E\in\mathbb{R}^{(m+1)\times d_{\mathrm{model}}},$
containing the $m$ item embeddings plus $E[\textsc{null}]=W_{\mathrm{TE}}[\mathrm{EOS}]$, preventing the model from “hiding” structure in embeddings and keeping the bottleneck aligned with QK capacity.